# Retrieve Only When It Needs: Adaptive Retrieval Augmentation for Hallucination Mitigation in Large Language Models

### v2

@@ 4.2 Baseline Methods
We consider the following methods as our baselines: (1) Vanilla LLMs, such as ChatGPT. (2) Self-improvement methods: CoVe Dhuliawala et al. (2023) generates verification questions to self-analyze potential errors, systematically addressing each question to refine the baseline response. Self-Reflection Ji et al. (2023b) presents an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation. Multi-agent Debate Du et al. (2023) utilize multiple LM agents to debate their individual responses over multiple rounds to arrive at a common final answer. (3) Retrieval-augmented methods: Factool Chern et al. (2023) leverages various tools to gather evidence about the factuality of the generated content. Detect and Mitigate Varshney et al. (2023) actively detects hallucinations during generation by identifying potential hallucination through the logit output values of LLMs. (4) Adaptive retrieval methods: FLARE Jiang et al. (2023) adopts an active retrieval strategy that only retrieves when LLMs generate low probability tokens. Adaptive-Retrieval Mallen et al. (2023) only retrieves when necessary based on a pre-defined threshold for entity popularity. Self-RAG Asai et al. (2023) trains a single arbitrary LM that adaptively retrieves passages on-demand. Adaptive-RAG Jeong et al. (2024) trains a query-complexity classifier to decide when to retrieve based on question complexity. LUQ Zhang et al. (2024) is a sampling-based uncertainty quantification for long text.
4.3 Implementation Details
@@
In our experiments, we validate Rowen using the ChatGPT language model. We re-implement all baselines, except Self-RAG, using ChatGPT to ensure a fair comparison. For semantic perturbations, we configure the temperature to 1.0 to generate diverse perturbed expressions. Otherwise, the temperature is set to 0.0 to obtain high-quality deterministic outputs. Considering the diversity of expressions and the latency in generating perturbations, we produce semantically equivalent questions. For the cross-language detection module, English serves as the source language while Chinese is employed as the target language. For the cross-model detection module, we adopt Qwen-Max-0428333https://qwenlm.github.io/blog/qwen-max-0428, a large instruction-tuned model for chat service, as the verifier LM. Following Chern et al. (2023), we utilize the Google Search API offered by Serper444https://serper.dev/ to search the top pages and extract the most pertinent search snippets from the API’s response. We also conduct extensive experiments to study the impact of various hyper-parameters in Appendix D.
5 Experimental Results
5.1 Main Results on Hallucination Mitigation
@@
| Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 23.0 | 75.6 | 20.0 |
@@
5.2 Effect of Detection Module
To validate the superiority of our hallucination detection module, we compare its performance in adaptive retrieval scenarios for hallucination mitigation with strong detection methods: (1) average token-level probability / entropy that utilizes the probabilities / entropies of tokens generated by a proxy LM (e.g., LLaMa2-7B) as a metric to measure hallucination. (2) SelfCheckGPT Manakul et al. (2023) that measures information consistency between the different responses to determine hallucinations. (3) Consistency-based method Zhang et al. (2023a), SAC3, that evaluates semantic-aware cross-check consistency, building upon the foundation of self-consistency principles.
Based on the results in Table 2, it is evident that logits-based methods perform moderately well in detecting hallucinations, especially on the StrategyQA dataset. Specifically, the entropy-based estimation exhibits superior performance. Besides, the combination of different variants in SelfCheckGPT leads to a slight performance improvement. Additionally, it is worth noting that SAC3 achieve competitive performance on detecting potential factual errors accurately.
Finally, we observe that Rowen significantly outperforms these strong hallucination detection baselines, especially the monolingual detection method SAC3. We also notice that Rowen achieves notable hallucination mitigation with a minimal number of retrieval calls. This underlines the superior efficiency of our hallucination detection module.
5.3 Scalability of Rowen
@@
In addition to ChatGPT, we also assess Rowen’s effectiveness when employing open-source language models: Qwen1.5-14B-Chat, Qwen2-7B-Instruct, and Qwen2-72B-Instruct555https://huggingface.co/Qwen. These models are chosen for the cross-language detection model due to their strong capabilities in following Chinese instructions, a critical feature for effective cross-language detection. The Llama-series models are not considered due to their weaker performance in generating Chinese responses. Instead, the Qwen-series models have demonstrated state-of-the-art performance in Chinese, which aligns with our research goals. For verifier LM , we choose to use Llama-3-8B-Instruct666https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
@@
5.4 Analysis of Computation Cost for Rowen
Efficiency Analysis of LLM Calls
To evaluate the efficiency of LLM calls in Rowen, we compare various methods—including Self-Reflection, Multi-agent Debate and three different Rowen variants—by analyzing the number of API calls required to answer a single question and their effectiveness in mitigating hallucinations. From the results shown in Table 4, we find that Rowen variants, particularly Rowen-CL and Rowen-CM, achieve significantly better hallucination mitigation compared to the baseline methods, while using a similar number of API call. To further improve efficiency, We design two useful speedup strategies to improve the efficiency of Rowen and describe them in Appendix E.
@@ 5.4 Analysis of Computation Cost for Rowen
Efficiency Analysis of Retrieval Calls
To verify the retrieval efficiency of Rowen, in Table 5, we compare the average number of retrieval calls made by Factool, Detect-and-Mitigate, FLARE, and Rowen-Hybrid to answer a question across two datasets. Rowen-Hybrid, which retrieves information only for uncertain responses, excels in both retrieval efficiency and factual accuracy, showcasing its superiority over other RAG methods.
@@
5.5 Quantitative Analysis of Hallucinations
6 Conclusion and Future Work
@@
In this paper, we introduce Rowen, a novel method designed to actively identify potential hallucinated responses in LLMs and correcting factual errors through an adaptive retrieval process. Our experiments on four datasets demonstrate the effectiveness of our approach. We also conduct detailed analytical experiments to underscore the impact of each module and the choices of hyper-parameters. Moving forward, we plan to expand our work by exploring how to effectively utilize retrieved evidence, even when irrelevant documents are integrated, and devise an effective method to minimize the influence of knowledge conflicts between internal knowledge and retrieved documents.
Limitations
Our method has several drawbacks. Firstly, it focuses on post-hoc correction, limiting its ability to detect hallucination only after generating a complete sentence rather than in real-time during the generation of contexts. Secondly, for ensuring accurate answers, our approach requires multiple requests to the API interface. This involves generating the CoT answer, perturbing and responding to the input, correcting the original answer, and so on, leading to increased latency. To address this, we can expedite the entire process by parallelizing API calls for both the OpenAI API and Google Search API. Additionally, reducing the number of perturbations can strike a better balance between detection precision and inference speed. Thirdly, our approach neglects to analyze and utilize the factual nature of the retrieved evidence, which we leave for future work.
Ethics Statement
This study delves into the detection and mitigation of hallucinations in language models (LMs), presenting broader implications for the realm of natural language processing (NLP) and addressing ethical concerns related to trust and reliability. Our method has the potential to enhance the accuracy and dependability of LMs, minimizing the risks associated with misinformation and biased outputs. Moreover, it foster accountability and trust in AI systems, contributing to the ongoing development of more reliable language models.
Appendix A Kernel Density Estimation Visualization
In our work, we hypothesize that lower consistency levels correlate with increased uncertainty in the model, consequently leading to a higher probability of inaccurate predictions. To verify this hypothesis, we analyze the consistency distribution through kernel density estimation plots on the TruthfulQA dataset. As illustrated in Figure 5, in the lower range of consistency check scores (left-skewed), the proportion of incorrect predictions is higher. This indicates that low consistency check scores are often associated with incorrect model predictions. On the contrary, in the higher range of consistency check scores (right-skewed), the proportion of correct predictions is higher. This indicates that high consistency check scores are usually associated with correct model predictions. There is some overlap between the distributions of different labels, but the overall trend shows that the higher the consistency check score, the higher the likelihood of correct predictions. These findings suggest that consistency check scores can be an effective indicator of prediction accuracy: the higher the score, the more accurate the model prediction is likely to be. These findings provide empirical support for our hypothesis. Additionally, we observe that the suitable threshold points for distinguishing correct from incorrect answers differ between Rowen-CL and Rowen-CM, with the optimal threshold for Rowen-CM being higher. This observation aligns with the conclusion in § D.
Appendix B Related Work on Exploring Uncertainty for Hallucination Detection
Uncertainty refers to the confidence level of the model outputs, and it serves an important indicator for identifying and eliminating hallucinations, so it can assist users in determining when to trust LLMs. In general, methods for exploring uncertainty for hallucination detection can be categorized into three types: (1) Logit-based estimation relies on accessing the model’s logits to calculate token-level probabilities or entropy, which are used to measure uncertainty Huang et al. (2023b); Varshney et al. (2023). However, this approach can pose challenges for black-box closed-source models. (2) Verbalized-based estimation involves prompting language models to express their uncertainty using specific prompts Xiong et al. (2023); Tian et al. (2023); Agrawal et al. (2023). However, these methods tend to display a high degree of overconfidence when expressing their confidence Xiong et al. (2023); Tao et al. (2024). (3) To overcome these limitations, consistency-based estimations are adopted to measure the consistency score among multiple responses provided by the model for a given question Manakul et al. (2023); Xiong et al. (2023); Wang et al. (2023b); Zhao et al. (2023). The underlying assumption suggests that when language models struggle with indecision and fabricate facts, they tend to provide logically inconsistent responses to identical questions. In this work, we propose that cross-language and cross-model consistency can offer highly sensitive signals for identifying hallucinations. Therefore, we utilize cross-language and cross-model detection modules that cross-check answers to the same question across different languages or models. This cross-checking paradigm serves as a powerful mechanism to identify hallucinations in LLMs.
Appendix C Discussions
There is a line of recent studies that concentrate on mitigating hallucinations with adaptive retrieval Jiang et al. (2023); Mallen et al. (2023); Asai et al. (2023); Jeong et al. (2024). Our work diverges from existing methodologies by: (1) Traditional self-consistency method Wang et al. (2023b) may falter when LLMs provide coherent yet incorrect responses to the same query, which is observed in SAC3. Our approach extends self-consistency uncertainty estimation to multilingual and multi-model settings, yielding superior hallucination mitigation performance compared to conventional methods (see Table 1). (2) FLARE Jiang et al. (2023) relies on the token probability of model outputs, a limitation for commercial closed-source models. In contrast, our method overcomes this constraint by assessing semantic similarity in generated content, which is suitable for white / black-box models. (3) Self-RAG Asai et al. (2023) necessitates intricate retraining processes and significant computational resources (4 Nvidia A100 GPUs and 80GB memory). In contrast, our method effectively mitigates hallucination without the need for retraining, underscoring its suitability for resource-constrained scenarios. (4) Adaptive-Retrieval Mallen et al. (2023) employs Wikipedia entity popularity for retrieval decisions, which proves ineffective for adversarial questions lacking clear entities in datasets like TruthfulQA. (5) Adaptive-RAG Jeong et al. (2024) dynamically decide whether to retrieve based on a pre-trained classifier, which is a smaller LM trained to predict the complexity level of incoming queries. Adaptive-RAG could be influenced by the accuracy of the classifier, given the absence of training datasets for query-complexity classifier, potentially leading to incorrect labeling of queries.
Our approach, examining linguistic diversity (cross-language) and model response consistency (cross-model), effectively identifies and mitigates hallucinations. This enhances model robustness and reliability, providing a solid foundation for applying LLMs. Our method is novel, practical, flexible, training-free, and broadly applicable, offering new insights and optimizing LLM research and application.
Appendix D Impact of Hyper-parameters in Rowen
Impact of Detection Threshold.
We experiment with various consistency thresholds to investigate their impact on hallucination mitigation effectiveness. As shown in Figure 6, we observe significant performance improvement when the detection threshold increases from 0.2 to 0.6, as this identifies more hallucinations and enhances factual accuracy by retrieving more external information. However, further increasing the threshold introduces more noise and irrelevant evidence, degrading performance. This suggests that an appropriate threshold is crucial for balancing detection accuracy and the retrieval process. Additionally, Rowen-CL and Rowen-CM have different optimal thresholds of 0.6 and 0.8, respectively, and the Rowen-Hybrid method demonstrates the most stable performance across varying thresholds.
Impact of Number of Perturbed Questions.
We investigate the impact of the number of perturbed questions of Rowen on the performance of hallucination mitigation. The results in Figure 7 reveal that the mitigation performance improves with the increase in the number of perturbations, particularly notable in the transition from 4 to 6 perturbations. This enhancement is attributed to the greater number of perturbations aiding in higher detection accuracy. However, further increasing the perturbation count beyond 6 yields only marginal improvements in performance (even no improvements). Thus, after considering the results, we determine that the optimal number of perturbations for achieving the best hallucination mitigation performance is 6.
Impact of Cross-Model Detection Weight.
We also examine the impact of different cross-model detection weight for Rowen-Hybrid. We ensure that for any given , the detection threshold is recalibrated to maintain optimal performance. Figure 8 shows that GPT-Judge scores significantly increase as the weight rises from 0.25 to 1.0, indicating that introducing cross-model checking can improve the accuracy of hallucination detection. However, when the weight factor exceeds 1.0, performance declines, with scores dropping to around 58 at a weight of 1.5. This suggests that while a moderate weight enhances factuality, excessively a high value allows the cross-model detection module to exert too much influence, reducing effectiveness.
Impact of the Choice of Target Language
| Language Pair | Discrepancy | GPT-Judge |
| English - German | Low | 50.55 |
| English - French | Medium | 52.26 |
| English - Chinese | High | 57.39 |
Impact of the Choice of Verifier LM
To investigate the impact of different verifier LMs on the efficacy of hallucination mitigation for Rowen-CM, we conducted experiments using various verifier LMs on the TruthfulQA dataset. The experimental results presented in Table 7 demonstrate a significant improvement in hallucination mitigation as the capability of the verifier LM increases. Specifically, Qwen-Max exhibited the highest effectiveness in this task, achieving an efficacy of 56.29%, while Qwen-Turbo and Qwen-Plus achieved 52.50% and 54.09%, respectively. These findings underscore the critical role of verifier LM selection in enhancing the effectiveness of hallucination mitigation.
| Verifier LM | Capability | GPT-Judge |
| Qwen-Turbo | Low | 52.50 |
| Qwen-Plus | Medium | 54.09 |
| Qwen-Max | High | 56.29 |
Impact of Perturbation Objects
In evaluating cross-language and cross-model perturbations, as well as calculating subsequent consistency scores, applying CL and CM perturbations to the original question appears to be a viable approach. However, recent studies suggest that merely perturbing questions might lead to responses that are consistent but not necessarily accurate Zhang et al. (2023a). This implies that question-level consistency alone does not suffice for assessing factual correctness, as these methods predominantly emphasize semantic coherence.
To validate this, we conduct preliminary experiments by applying CL and CM perturbations to the original question (question-level consistency checking) and compare the results with those of Rowen. The results, as shown in Table 8, demonstrate that Rowen consistently outperforms the approach of applying perturbations to the original question. This reinforces that Rowen remains the most effective method for ensuring accuracy.
| Object | Variants | TruthfulQA | StrategyQA |
| GPT-Judge | Accuracy | ||
| Input | CL | 55.34 | 72.00 |
| CM | 55.85 | 72.00 | |
| Perturbations | CL | 57.39 | 74.00 |
| CM | 56.29 | 72.40 |
Appendix E Strategies to improve efficiency
We have implemented two helpful strategies to improve the efficiency without compromising the performance by:
(1) Our cross-language detection and cross-model detection modules can operate in parallel. This significantly reduces latency, as it allows for the simultaneous evaluation of semantic consistency across different languages and models.
(2) We have optimized the generation process for responding to multiple semantically equivalent perturbed questions by leveraging parallel API calls. This approach expedites the retrieval of diverse responses, which are crucial for our consistency checks.
By employing these parallel processing strategies, we ensure that the efficiency of our method is improved while preserving its effectiveness. We believe these optimizations adequately address the concerns about the method’s efficiency and demonstrate our commitment to developing a scalable and high-performing framework.
Appendix F Prompt Instruction Examples
In this section, we provide these instructions used in our experiments.
| Transform the following question into a query suitable for information retrieval and then use the query to find relevant evidence. Follow these steps for a short but comprehensive approach: |
| Original Question: |
| 1. Simplify the question into key terms and concepts. |
| 2. Remove any elements that are not essential for a search query. |
| 3. If necessary, rephrase the question to focus on the most critical aspects for retrieval. |
| You are an advanced AI with the ability to process and synthesize information efficiently. When provided with a question, you should consider the following aspects in your response: |
| 1. The specific question asked: |
| 2. The reasoning process: |
| 3. The short answer previously given related to the question: |
| 4. Evidences that have been retrieved: |
| Use this information to generate a concise, accurate, and relevant answer that reflects the latest understanding of the topic based on the input provided. |
| Your response should be clear, direct, and provide the most up-to-date information available while maintaining coherence with the previous discussion points. |

### v3

@@ 4.2. Baseline Methods
We consider the following methods as our baselines: (1) Vanilla LLMs, such as ChatGPT. (2) Self-improvement methods: CoVe (Dhuliawala et al., 2023) generates verification questions to self-analyze potential errors, systematically addressing each question to refine the baseline response. Self-Reflection (Ji et al., 2023b) presents an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation. Multi-agent Debate (Du et al., 2023) utilize multiple LM agents to debate their individual responses over multiple rounds to arrive at a common final answer. (3) Retrieval-augmented methods: Factool (Chern et al., 2023) leverages various tools to gather evidence about the factuality of the generated content. Detect and Mitigate (Varshney et al., 2023) actively detects hallucinations during generation by identifying potential hallucination through the logit output values of LLMs. (4) Adaptive retrieval methods: FLARE (Jiang et al., 2023) adopts an active retrieval strategy that only retrieves when LLMs generate low probability tokens. Adaptive-Retrieval (Mallen et al., 2023) only retrieves when necessary based on a pre-defined threshold for entity popularity. Self-RAG (Asai et al., 2023) trains a single arbitrary LM that adaptively retrieves passages on-demand. Adaptive-RAG (Jeong et al., 2024) trains a query-complexity classifier to decide when to retrieve based on question complexity. LUQ (Zhang et al., 2024) is a sampling-based uncertainty quantification for long text.
4.3. Implementation Details
LLMs
@@ 4.3. Implementation Details
In our experiments, we validate Rowen using the GPT-3.5 language model, specifically the gpt-3.5-turbo-0613 version333https://api.openai.com/v1/chat/completions. We re-implement all baselines, except Self-RAG, using GPT-3.5 to ensure a fair comparison. For semantic perturbations, we configure the temperature to 1.0 to generate diverse perturbed expressions. Otherwise, the temperature is set to 0.0 to obtain high-quality deterministic outputs. Considering the diversity of expressions and the latency in generating perturbations, we produce semantically equivalent questions. For the cross-language detection module, English serves as the source language while Chinese is employed as the target language444We study the impact of the choice of target language in § 5.4 Table 7.. This decision is made considering the substantial cultural disparities between the two languages, which can enhance the model’s capability to detect semantic inconsistencies when responding to identical questions. For the cross-model detection module, we adopt Qwen-Max-0428555https://qwenlm.github.io/blog/qwen-max-0428, a large instruction-tuned model for chat service, as the verifier LM666We also study the impact of the choice of verifier LM in § 5.4 Table 8.. We execute all experiments on 8 NVIDIA A800 80G GPUs.
@@
| Rowen-Hybrid | 59.34 | 15.27 | 31.15 | 23.0 | 75.60 | 20.0 |
Retrieval Module
Following Chern et al. (2023), we utilize the Google Search API offered by Serper777https://serper.dev/ to search the top pages and extract the most pertinent search snippets from the API’s response. Subsequently, we parse the response to acquire various types of snippets, including answer boxes, knowledge graphs, and organic search results.
5. Experimental Results
In this section, we evaluate the effectiveness of Rowen for hallucination mitigation. In particular, we aim to answer the following research questions:
RQ1: How does Rowen perform compared to existing strong methods for hallucination mitigation?
RQ2: Is the proposed multilingual detection module superior to other hallucination detection methods?
RQ3: How well does Rowen generalize to new datasets and models?
RQ4: How do different hyperparameters affect the performance of Rowen?
RQ5: Does Rowen eliminate both types of hallucinations as expected?
RQ6: Does Rowen reduce unnecessary retrieval calls?
5.1. Main Results (RQ1)
@@
5.2. Effect of Detection Module (RQ2)
To validate the effectiveness of our proposed hallucination detection module, we compare its performance in adaptive retrieval scenarios for hallucination mitigation with strong detection methods: (1) average token-level probability / entropy that utilizes the probabilities / entropies of tokens generated by a proxy LM (e.g., LLaMa2-7B) as a metric to measure hallucination. (2) SelfCheckGPT (Manakul et al., 2023) that measures information consistency between the different responses to determine hallucinations. (3) Consistency-based method (Zhang et al., 2023c), SAC3, that evaluates semantic-aware cross-check consistency, building upon the foundation of self-consistency principles.
Based on the results in Table 5, it is evident that logits-based methods perform moderately well in detecting hallucinations. Specifically, the entropy-based estimation exhibits superior performance. Besides, the combination of different variants in SelfCheckGPT leads to a slight performance improvement. Additionally, it is worth noting that SAC3 achieve competitive performance on detecting potential factual errors accurately.
Finally, we observe that Rowen significantly outperforms these strong hallucination detection baselines, especially the monolingual detection method SAC3. We also notice that Rowen achieves notable hallucination mitigation with a minimal number of retrieval calls. This underlines the superior efficiency of our adaptive retrieval module.
5.3. Scalability of Rowen (RQ3)
@@
In addition to ChatGPT, we also assess Rowen’s effectiveness when employing open-source language models: Qwen1.5-14B-Chat, Qwen2-7B-Instruct, and Qwen2-72B-Instruct888https://huggingface.co/Qwen. These models are chosen for the cross-language detection model due to their strong capabilities in following Chinese instructions, a critical feature for effective cross-language detection. The Llama-series models are not considered due to their weaker performance in generating Chinese responses. Instead, the Qwen-series models have demonstrated state-of-the-art performance in Chinese, which aligns with our research goals. For verifier LM , we choose to use Llama-3-8B-Instruct999https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
@@
5.4. Impact of Hyper-parameters in Rowen (RQ4)
We analyze the impact of key factors for hallucination mitigation on the TruthfulQA dataset, including detection thresholds, numbers of perturbed questions, and cross-model detection weights, as shown in Figure 4.
Detection Threshold
We experiment with various consistency thresholds to investigate their impact on hallucination mitigation effectiveness. As shown in Figure 4(a), we observe significant performance improvement when the detection threshold increases from 0.2 to 0.6, as this identifies more hallucinations and enhances factual accuracy by retrieving more external information. However, further increasing the threshold introduces more noise and irrelevant evidence, degrading performance. This suggests that an appropriate threshold is crucial for balancing detection accuracy and the retrieval process. Additionally, Rowen-CL and Rowen-CM have different optimal thresholds of 0.6 and 0.8, respectively, and the Rowen-Hybrid method demonstrates the most stable performance across varying thresholds.
Number of Perturbed Questions
We also study the effect of the number of perturbed question on hallucination mitigation and the results are shown in Figure 4(b). We observe that the performance of Rowen improves with an increasing number of question samples, yet the performance gain gradually diminishes after surpassing 6 question samples. This indicates that, in practice, using 6 question samples could achieve reasonably good performance at a relatively low computational cost.
Cross-Model Detection Weight
We also examine the impact of different cross-model detection weight for Rowen-Hybrid. Figure 4(c) shows that GPT-Judge scores significantly increase as the weight rises from 0.25 to 1.0, indicating that introducing cross-model checking can improve the accuracy of hallucination detection. However, when the weight factor exceeds 1.0, performance declines, with scores dropping to around 56 at a weight of 1.5. This suggests that while a moderate weight enhances factuality, excessively a high value allows the cross-model detection module to exert too much influence, reducing effectiveness.
Impact of the Choice of Target Language
| Language Pair | Discrepancy | GPT-Judge |
| English - German | Low | 50.55 |
| English - French | Medium | 52.26 |
| English - Chinese | High | 57.39 |
| Verifier LM | Capability | GPT-Judge |
| Qwen-Turbo | Low | 52.50 |
| Qwen-Plus | Medium | 54.09 |
| Qwen-Max | High | 56.29 |
Impact of the Choice of Verifier LM
To investigate the impact of different verifier LMs on the efficacy of hallucination mitigation for Rowen-CM, we conducted experiments using various verifier LMs on the TruthfulQA dataset. The experimental results presented in Table 8 demonstrate a significant improvement in hallucination mitigation as the capability of the verifier LM increases. Specifically, Qwen-Max exhibited the highest effectiveness in this task, achieving an efficacy of 56.29%, while Qwen-Turbo and Qwen-Plus achieved 52.50% and 54.09%, respectively. These findings underscore the critical role of verifier LM selection in enhancing the effectiveness of hallucination mitigation tasks.
5.5. Quantitative Analysis (RQ5)
@@
5.6. Analysis of Computation Cost (RQ6)
Efficiency Analysis of LLM Calls
To evaluate the efficiency of LLM calls in Rowen, we compare various methods—including Self-Reflection, Multi-agent Debate and three different Rowen variants—by analyzing the number of API calls required to answer a single question and their effectiveness in mitigating hallucinations. From the results shown in Table 9, we find that Rowen variants, particularly Rowen-CL and Rowen-CM, achieve significantly better hallucination mitigation compared to the baseline methods, while using a similar number of API call.
Efficiency Analysis of Retrieval Calls
To verify the retrieval efficiency of Rowen, in Table 10, we compare the average number of retrieval calls made by Factool, Detect-and-Mitigate, FLARE, and Rowen-Hybrid to answer a question across two datasets. Factool generates the most retrieval API calls due to verifying each claim. Detect-and-Mitigate and FLARE identify low-confident concepts in LLM outputs and call search APIs with fewer API calls. Adaptive-RAG underestimates the difficulty of adversarial questions in the TruthfulQA dataset, resulting in poor truthfulness scores despite using the minimum number of retrievals. Rowen-Hybrid, which retrieves information only for uncertain responses, excels in both retrieval efficiency and factual accuracy, showcasing its superiority over other RAG methods. This demonstrates the superiority of Rowen-Hybrid over other RAG methods in terms of both efficiency and effectiveness.
6. Conclusion and Future Work
@@
In this paper, we introduce Rowen, a novel method designed to actively identify potential hallucinated responses in LLMs and correcting factual errors through an adaptive retrieval process. Our experiments on two datasets demonstrate the effectiveness of our approach. We also conduct detailed analytical experiments to underscore the impact of each module and the choices of hyper-parameters. Moving forward, we plan to expand our work by exploring how to effectively utilize retrieved evidence, even when irrelevant documents are integrated, and devise an effective method to minimize the influence of knowledge conflicts between internal knowledge and retrieved documents.
Acknowledgement
This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No. 62276248, the Key Research and Development Program of Xinjiang Uyghur Autonomous Region Grant No. 2024B03026, the Beijing Nova Program under Grants No. 20250484765, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.